OpenStudy (anonymous):

Traveling against the current a ship went 220 km in 3 hours & 20 minutes. The return trip with the current took 35 minutes less. Find (a) the ship's rate in still water and (b) the rate of the current. Please explain.

OpenStudy (sirm3d):

let s = ship's rate in still water c = rate of the current effective rate against the current is \(s-c\) effective rate with the current is \(s+c\) distance=rate x time against the current: \(220=(s-c)\times 200 \text{ (minutes)}\) with the current: \(220=(s+c)\times 165 \text{ (35 minutes shorter/faster)}\)